OpenStudy (anonymous):

My teacher on FLVS is on vacation. Would somebody please help explain this for me? Thank you! (Solve for x: −4|x + 5| = −16)

OpenStudy (anonymous):

first divide -4 on both sides, that will get you |x+5|=4

OpenStudy (anonymous):

So, x+5=4... So x=-1?

OpenStudy (anonymous):

"||" these things mean determinant. so there are two possibilities.

OpenStudy (anonymous):

first take these out "||". and write that there are two possibilities .

OpenStudy (anonymous):

first one is x+5 = -4 and the other one is x+5 = +4

OpenStudy (anonymous):

So, the possibilities are -9 and -1?

OpenStudy (anonymous):

x+5= -4 => x=-9 and x+5 = 4 => x= -1

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ah, thank you soo much!

OpenStudy (anonymous):

then write answer like this, "solution set = {-9, -1}". this is the correct way to write this kind of answer.

OpenStudy (anonymous):

no problem